OpenStudy (anonymous):

Will fan and medal! 2. (10)9-(10)11 a. -2 b. -20 c. 200 d. 8 4. 6(6)+6(4) a. 60 b. 10 c. 16 d. 12 6. Use the distributive property to multiply: 5(2t-5) a.10t-5 b. 7t-25 c. 10t-25 d. -15t

OpenStudy (zzr0ck3r):

9*10-10*11 = 90-110 = ?

OpenStudy (anonymous):

I dont understand

OpenStudy (zzr0ck3r):

do you understand 9(10) = 9*10 = 90?

OpenStudy (anonymous):

Sort of. But what is *? Does it mean divid?

OpenStudy (anonymous):

How would 9 divided by 10 equal 90? 9 times 10 equals 90

OpenStudy (anonymous):

Thats why I was confused. Sorry to inconvenience you. @FreeWilly.